OpenStudy (anonymous):

Point K is graphed on a number line at 3. Point L is 11 units away from point K on the number line. What could be the coordinates of point L? Choose all answers that are correct. A. –14 B. –8 C. 7 D. 14 (Multiple Answers)

OpenStudy (texaschic101):

ok...your at 3...and you have to move 11 spaces...in both directions 3 + 11 = 14 3 + (-11) = 3 - 11 = - 8 so those are your 2 points that are 11 spaces from 3 any questions ?

OpenStudy (anonymous):

so what is the answers?

OpenStudy (texaschic101):

14 and -8 are your answers

OpenStudy (anonymous):

ty

OpenStudy (texaschic101):

YW :)